# Omni-Recon: Towards General-Purpose Neural Radiance Fields for Versatile 3D Applications

### v1

@@ Abstract
Recent breakthroughs in Neural Radiance Fields (NeRFs) have sparked significant demand for their integration into real-world 3D applications. However, the varied functionalities required by different 3D applications often necessitate diverse NeRF models with various pipelines, leading to tedious NeRF training for each target task and cumbersome trial-and-error experiments. Drawing inspiration from the generalization capability and adaptability of emerging foundation models, our work aims to develop one general-purpose NeRF for handling diverse 3D tasks. We achieve this by proposing a framework called Omni-Recon, which is capable of (1) generalizable 3D reconstruction and zero-shot multitask scene understanding, and (2) adaptability to diverse downstream 3D applications such as real-time rendering and scene editing. Our key insight is that an image-based rendering pipeline, with accurate geometry and appearance estimation, can lift 2D image features into their 3D counterparts, thus extending widely explored 2D tasks to the 3D world in a generalizable manner. Specifically, our Omni-Recon features a general-purpose NeRF model using image-based rendering with two decoupled branches: one complex transformer-based branch that progressively fuses geometry and appearance features for accurate geometry estimation, and one lightweight branch for predicting blending weights of source views. This design achieves state-of-the-art (SOTA) generalizable 3D surface reconstruction quality with blending weights reusable across diverse tasks for zero-shot multitask scene understanding. In addition, it can enable real-time rendering after baking the complex geometry branch into meshes, swift adaptation to achieve SOTA generalizable 3D understanding performance, and seamless integration with 2D diffusion models for text-guided 3D editing. All code will be released upon acceptance.
@@ 1 Introduction
  • Extensive experiments demonstrate that our Omni-Recon can achieve state-of-the-art (SOTA) generalizable 3D surface reconstruction quality and SOTA generalizable 3D understanding performance. Additionally, it can support fast reconstruction and real-time rendering of a new scene, along with easy-to-use 3D scene editing.
Our Omni-Recon underscores the potentially wide usage of image-based rendering pipelines in diverse real-world 3D applications, which may have been overlooked by previous works. We believe that the insights we provide could ignite future innovations in more advanced rendering pipelines.
@@ 2 Related Work
Efficient scene representations. Various scene representations have been proposed to accelerate NeRFs’ training and inference process. To speed up NeRF training, [62, 20, 9, 43, 8] integrate explicit neural representations with volumetric rendering. In parallel, to speed up NeRF inference/rendering,  [35, 52, 53] attempt to reduce the complexity of the MLP model in NeRF. In addition, [22, 87, 25, 27, 37, 59, 53] explore the free space in a target 3D scene to improve the sampling efficiency and thus reduce unnecessary MLP inference. Motivated by the insight of deferred rendering [67][25, 12, 85, 65] pre-compute scene features offline and thus require less computation at run-time. In particular, [12, 85, 65] represents NeRF as a set of textured polygons, where the textures can be pre-computed and thus the NeRF pipeline can be efficiently executed in modern graphics hardware. However, these efficient scene representations still necessitate costly per-scene optimization, lacking generalization capabilities.
Figure 1: An overview of our Omni-Recon framework.
@@ 3.2 The Overall Image-Based Rendering Pipeline
$s=\mathbf{M}_{sdf}(\{\mathbf{f}_{i}^{+}\}_{i=1}^{N},V),\,\,\,\,\,\,\{\omega_{i}\}_{i=1}^{N}=\mathbf{M}_{color}(\{\mathbf{f}_{i}^{+}\}_{i=1}^{N},\mathbf{d}),$ (1)
@@ 3.3 The Proposed Transformer-based Geometry Branch
Overview. Our geometry branch leverages transformer modules [71] to progressively fuse geometry and appearance features, thus properly handling the two occlusion effects mentioned above. Specifically, for the $K$ sampled 3D points along the ray, we acquire their geometry features $\{\mathbf{v}_{k}\}_{k=1}^{K}$ from the geometry feature volume $V$, which are sequentially processed through $B$ blocks ($B$=2 in our design), each consisting of a geometry transformer, an appearance transformer, and an occlusion transformer, to produce the per-point SDF. Note that we leverage geometry feature volumes as inputs to the first block since they carry the richest geometric clues compared to the appearance features, as demonstrated in the supplementary materials.
@@
where $\mathbf{x}$ is the input to the geometry transformer, which is exactly the geometry features themselves if the geometry transformer in the first block, and $\mathbf{q}$, $\mathbf{k}$, and $\mathbf{v}$ are the query, key, and value of the attention scheme [71], respectively.
@@ 3.4 Training Objectives
Omni-Recon (Ours) 1.10 0.95 2.17 1.52 0.89 1.09 1.47 0.83 1.29 1.20 0.83 1.03 0.73 0.51 0.97 1.05
@@ 3.5 Evaluation: Generalizable 3D Reconstruction
Sparse view reconstruction. We benchmark the performance of sparse view reconstruction using three views from each testing scene in DTU, following the setting in [16, 55, 34]. As shown in Tab. 1, we can observe that (1) Our Omni-Recon can achieve new SOTA reconstruction performance, with the lowest Chamfer distance averaged over all scenes. Additionally, as visualized in Fig. 2 (Row 1), our reconstructed mesh can exhibit smooth surfaces and correctly maintain the structure, e.g., the hole in the mouth. (2) Our method achieves the best reconstruction performance in 11 out of 15 scenes and the second-best in the remaining 4 scenes, indicating its general effectiveness across diverse scenes.
@@ 4 Omni-Recon: Enable Zero-shot Scene Understanding
To enable the understanding of various 3D scene properties, one approach is to jointly train the NeRF backbone on the target scene understanding tasks. However, this approach is not scalable due to the diversity of scene properties, such as semantic segmentation, edge detection, and keypoint prediction, making it difficult to cover all of them. Therefore, it is highly desirable to enable zero-shot scene understanding for new tasks that were not seen during training.
@@ 5.4 Evaluation: Effectiveness on Different Downstream 3D Tasks
Observations and analysis: As shown in Sec. 4, we can observe that (1) Without any finetuning, the rendering pipeline already exhibits certain rendering capabilities, despite the model backbone being trained using volumetric rendering with more than one sampled point; (2) Our rendering pipeline can be swiftly finetuned to significantly boost the PSNR. For instance, with just a 10-second finetuning, our rendering pipeline can match the average rendering quality of the strongest baseline, ReTR [34]; With 1-minute/2-minute finetuning, it can surpass ReTR by a +2.75 and +3.36 PSNR improvement on average, respectively; (3) Our rendering pipeline enables real-time rendering with 71.3 FPS, i.e., a $>$2458× speed-up over the baselines. Here, the assumption is that the feature extraction of each source view from a target scene can be done in a one-time effort and does not account for throughput. If we still consider this feature extraction latency for each rendering, the resulting 40.82 FPS still meets the real-time requirement.

### v3

# Omni-Recon: Harnessing Image-based Rendering for General-Purpose Neural Radiance Fields

@@ Abstract
Recent breakthroughs in Neural Radiance Fields (NeRFs) have sparked significant demand for their integration into real-world 3D applications. However, the varied functionalities required by different 3D applications often necessitate diverse NeRF models with various pipelines, leading to tedious NeRF training for each target task and cumbersome trial-and-error experiments. Drawing inspiration from the generalization capability and adaptability of emerging foundation models, our work aims to develop one general-purpose NeRF for handling diverse 3D tasks. We achieve this by proposing a framework called Omni-Recon, which is capable of (1) generalizable 3D reconstruction and zero-shot multitask scene understanding, and (2) adaptability to diverse downstream 3D applications such as real-time rendering and scene editing. Our key insight is that an image-based rendering pipeline, with accurate geometry and appearance estimation, can lift 2D image features into their 3D counterparts, thus extending widely explored 2D tasks to the 3D world in a generalizable manner. Specifically, our Omni-Recon features a general-purpose NeRF model using image-based rendering with two decoupled branches: one complex transformer-based branch that progressively fuses geometry and appearance features for accurate geometry estimation, and one lightweight branch for predicting blending weights of source views. This design achieves state-of-the-art (SOTA) generalizable 3D surface reconstruction quality with blending weights reusable across diverse tasks for zero-shot multitask scene understanding. In addition, it can enable real-time rendering after baking the complex geometry branch into meshes, swift adaptation to achieve SOTA generalizable 3D understanding performance, and seamless integration with 2D diffusion models for text-guided 3D editing. Our code is available at: https://github.com/GATECH-EIC/Omni-Recon.
@@ 1 Introduction
  • Extensive experiments demonstrate that our Omni-Recon can achieve state-of-the-art (SOTA) generalizable 3D surface reconstruction and scene understanding performance. Additionally, it can support real-time rendering of a new scene after rapid adaptation, along with easy-to-use 3D scene editing.
Our Omni-Recon underscores the potentially wide usage of image-based rendering pipelines in diverse real-world 3D applications, which may have been overlooked by previous works. We believe that the insights we provide could ignite future innovations in more advanced generalizable rendering pipelines.
@@ 2 Related Work
Efficient scene representations. Various scene representations have been proposed to accelerate NeRFs’ training and inference process. To speed up NeRF training, [62, 20, 9, 43, 8] integrate explicit neural representations with volumetric rendering. In parallel, to speed up NeRF inference/rendering,  [35, 52, 53] attempt to reduce the complexity of the multi-layer perceptron (MLP) model in NeRF. In addition, [22, 87, 25, 27, 37, 59, 53] explore the free space in a target 3D scene to improve the sampling efficiency and thus reduce unnecessary MLP inference. Motivated by the insight of deferred rendering [67][25, 12, 85, 65] pre-compute scene features offline and thus require less computation at run-time. In particular, [12, 85, 65] represents NeRF as a set of textured polygons, where the textures can be pre-computed and thus the NeRF pipeline can be efficiently executed in modern graphics hardware. However, these efficient scene representations still necessitate costly per-scene optimization, lacking generalization capabilities.
@@ 3.2 The Overall Image-Based Rendering Pipeline
$s=\mathbf{M}_{sdf}(\{\mathbf{f}_{i}\}_{i=1}^{N},V),\,\,\,\,\,\,\{\omega_{i}\}_{i=1}^{N}=\mathbf{M}_{color}(\{\mathbf{f}_{i}\}_{i=1}^{N},\mathbf{d}),$ (1)
@@ 3.3 The Proposed Transformer-based Geometry Branch
Overview. Our geometry branch leverages transformer modules [71] to progressively fuse geometry and appearance features, thus properly handling the two occlusion effects mentioned above. Specifically, for the $K$ sampled 3D points along the ray, we acquire their geometry features $\{\mathbf{v}_{k}\}_{k=1}^{K}$ from the geometry feature volume $V$, which are sequentially processed through $B$ blocks ($B$=2 in our design), each consisting of a geometry transformer, an appearance transformer, and an occlusion transformer, to produce the per-point SDF.
@@
where $\mathbf{x}$ is the input to the geometry transformer, which is exactly the geometry features themselves for the geometry transformer in the first block, and $\mathbf{q}$, $\mathbf{k}$, and $\mathbf{v}$ are the query, key, and value of the attention scheme [71], respectively.
@@ 3.4 Training Objectives
Omni-Recon (Ours) 1.13 0.91 2.13 1.52 0.93 1.09 1.70 0.84 1.29 1.20 0.83 1.04 0.81 0.55 1.05 1.05
@@ 3.5 Evaluation: Generalizable 3D Reconstruction
Sparse view reconstruction. We benchmark the performance of sparse view reconstruction using three views from each testing scene in DTU, following the setting in [16, 55, 34]. As shown in Tab. 1, we can observe that (1) Our Omni-Recon can achieve new SOTA reconstruction performance, with the lowest Chamfer distance averaged over all scenes. Additionally, as visualized in Fig. 2 (Row 1), our reconstructed mesh can exhibit smooth surfaces and correctly maintain the structure, e.g., the hole in the mouth. (2) Our method achieves the best reconstruction performance in 10 out of 15 scenes, indicating its general effectiveness across diverse scenes.
@@ 4 Omni-Recon: Enable Zero-shot Scene Understanding
To enable the understanding of various 3D scene properties, one approach is to jointly train the NeRF backbone on all target tasks. However, this approach is not scalable due to the diversity of scene properties, such as semantic segmentation, edge detection, and keypoint prediction, making it costly to cover new properties of interest. Therefore, it is highly desirable to enable zero-shot scene understanding for new tasks that were not seen during training.
@@ 5.4 Evaluation: Effectiveness on Different Downstream 3D Tasks
Observations and analysis: As shown in Tab. 4, we can observe that (1) Without any finetuning, the rendering pipeline already exhibits certain rendering capabilities, despite the model backbone being trained using volumetric rendering with more than one sampled point; (2) Our rendering pipeline can be swiftly finetuned to significantly boost the PSNR. For instance, with just a 10-second finetuning, our rendering pipeline can match the average rendering quality of the strongest baseline, ReTR [34]; With 1-minute/2-minute finetuning, it can surpass ReTR by a +2.75 and +3.36 PSNR improvement on average, respectively; (3) Our rendering pipeline enables real-time rendering with 71.3 FPS, i.e., a $>$2458× speed-up over the baselines. Here, the assumption is that the feature extraction of each source view from a target scene can be done in a one-time effort and does not account for throughput. If we still consider this feature extraction latency for each rendering, the resulting 40.82 FPS still meets the real-time requirement.
@@ 6 Conclusion
Acknowledgement
The work is supported by the National Science Foundation (NSF) through the SCH program (Award number: 1838873) and CoCoSys, one of the seven centers in JUMP 2.0, a Semiconductor Research Corporation (SRC) program sponsored by DARPA.